OpenStudy (anonymous):

slove for the unkown: -3(5X-4)=-21 and -18(y+.6)=27 also 6x-2y=13

OpenStudy (aroub):

-15x+12=-21 -15x=-21-12 -15x=-33 x=11/5

OpenStudy (aroub):

-18y+6=27 -18y=27-6 -18y=21 y=-7/6